# On the Effectiveness of Proposed Techniques to Reduce Energy Consumption in RAG Systems: A Controlled Experiment

### v1

@@ 3.4. Dataset
Several criteria needed to be considered for selecting a benchmark dataset. Firstly, the dataset had to be manageable, as very large-scale datasets were not suitable for our cluster. For example, the KILT benchmark (petroni2021kiltbenchmarkknowledgeintensive), designed for intensive language tasks, contains raw knowledge sources of approximately 35 GB, which would not be suitable for our experiment. Secondly, the dataset had to be rich and cover a diverse set of topics, as dataset diversity can make the results more convincing. Based on these criteria, we selected the CRAG dataset (crag_benchmark), which contains a manageable total of 2,706 questions. It also includes a diverse range of question types, such as comparison and multi-hop questions, which are reasonably close to real-world challenges commonly encountered in practical applications. CRAG highlights the inherent trade-offs between accuracy and latency across various state-of-the-art RAG systems, offering useful context for evaluating ChatRAG’s performance. The CRAG benchmark employs a structured scoring method to assess the quality of responses generated by RAG systems, categorizing each answer as perfect, missing, or incorrect. Following this approach, we implemented a similar labeling process. In our evaluation, we relied on an LLM-as-Judge approach (gu2025surveyllmasajudge) to validate the correctness of the final answers, excluding responses such as “I don’t know”. LLM-as-Judge approaches provide efficient ways to evaluate query responses that go beyond simple requests like multiple-choice questions and have proven remarkably effective for many benchmarks, matching even crowdsourced human evaluations (Zheng2023). We employed DeepSeek-V2 (deepseekai2024deepseekv2strongeconomicalefficient) as our LLM judge due to its low cost and competent question-answering capabilities. The details of this evaluation and the used prompts are available in our replication package [^13] https://anonymous.4open.science/r/green-rag-techniques-experiment/running/README.md.
@@ 3.6. Data Analysis
After we collected all the data from the experiments, we assessed the normality of each data subset using the Shapiro-Wilk test (Hanusz_Tarasinska_Zielinski_2016). Since all datasets in our experiment passed the normality test, we used the t-test (student1908probable) to evaluate the significance levels. Based on our hypotheses, we used a one-tailed test for energy consumption and a two-tailed test for performance. Since multiple hypotheses were tested simultaneously, we had to guard against the multiple comparisons problem, where the probability of incorrectly rejecting at least one true null hypothesis (Type I error) increases with the number of tests performed (BARNETT20222331). To mitigate this, we applied the Holm-Bonferroni correction (holm1979simple), which adjusts the p-values to control the family-wise error rate. After correction, an adjusted p-value ¡ 0.05 indicates a statistically significant difference, while a value ¿ 0.05 suggests no significant difference. For statistically significant differences, we calculated Cohen’s d (cohen2013statistical) to estimate the effect size. According to Cohen’s guidelines, values between 0.0 and 0.2 indicate a negligible effect, 0.2 to 0.5 a small effect, 0.5 to 0.8 a moderate effect, and 0.8 and above a large effect. Additionally, we calculated mean values for each configuration with a significant effect and then computed the percentage change relative to the baseline (control).
@@ 4.1. Reducing Energy Consumption (RQ1)
This limited effectiveness is likely due to the short prompt lengths and minimal overlap between retrieved documents across queries. To quantify query similarity, we applied TF-IDF vectorization (Rajaraman_Ullman_2011) combined with cosine similarity. The average similarity score was 0.0294 (standard deviation 0.0437), indicating that most queries differ substantially. As a result, the prefix cache was rarely reused, leading to negligible improvements. In contrast, systems with higher query similarity might benefit more from prefix caching.
@@ 4.3. Trade-off With Accuracy (RQ2b)
In this section, we first examine the baseline’s ability to answer questions accurately and then evaluate the impact of the proposed techniques on system accuracy. Among all 2,606 questions from our benchmark dataset, the baseline system produced, on average, correct answers for 619 questions, resulting in an accuracy of 23.75%. This is in line with the published CRAG benchmark results (crag_benchmark), which mention an accuracy of 23.7% for Llama 3.1 8B Instruct. Our baseline system achieved almost identical accuracy, which validates the realism and effectiveness of our base setup, but also of our LLM-as-Judge approach.
@@ 4.4. Combining the Two Best Techniques
Two technique configurations (T1 with the 0.78 threshold and T3 with 384 dimensions) emerged as the most beneficial candidates, improving energy consumption, latency, and even slightly accuracy. To understand whether their benefits compound when applied together, we ran an additional experiment with a system variation that implemented both of them simultaneously. As shown in Fig. 7, the combination of the two led to even greater improvements in energy consumption and latency. Specifically, T1-0.78 alone achieved a 20.7% reduction in energy consumption, while T3-384 alone yielded a 38.1% reduction. Their combination, however, reduced energy consumption by 46.7%. The results for latency were similarly beneficial. Although each individual configuration results in a modest accuracy improvement of 1.7%, their combination did not produce a statistically significant change in accuracy anymore. According to gu2021principled, as the dimensionality of embeddings increases, the distribution of pairwise cosine similarities tends to converge toward a stable distribution with finite variance. In higher-dimensional spaces, cosine similarity scores cluster more tightly around a central value, reducing the variance between similarity scores across different pairs. In practice, high-dimensional embeddings tend to make random vectors more orthogonal, concentrating pairwise cosine similarities near zero with low variance. This shifts the natural baseline similarity upward, e.g., from a slightly negative value in lower dimensions to near zero, which likely requires a threshold recalibration. Nonetheless, since the combination also did not harm accuracy, applying both techniques at once remains a powerful option to improve energy efficiency.
@@ 6. Threats to Validity
We assess potential threats to the validity of our experiment following the framework of Wohlin et al. (wohlin12).
Internal Validity. Chunking is critical for RAG systems, as larger chunks can improve recall but may reduce precision (pesl2025retrievalaugmentedgenerationservicediscovery). We used the TokenTextSplitter of Spring AI with a chunk size of 320 tokens to balance model compatibility and manageable chunk counts. Despite this, 10 out of 501,916 chunks failed, which could marginally affect accuracy if they contained relevant information.

### v2

@@ 3.4. Dataset
Several criteria needed to be considered for selecting a benchmark dataset. Firstly, the dataset had to be manageable, as very large-scale datasets were not suitable for our cluster. For example, the KILT benchmark (Petroni et al., 2021), designed for intensive language tasks, contains raw knowledge sources of approximately 35 GB, which would not be suitable for our experiment. Secondly, the dataset had to be rich and cover a diverse set of topics, as dataset diversity can make the results more convincing. Based on these criteria, we selected the CRAG dataset (Yang et al., 2024), which contains a manageable total of 2,706 questions. It also includes a diverse range of question types, such as comparison and multi-hop questions, which are reasonably close to real-world challenges commonly encountered in practical applications. CRAG highlights the inherent trade-offs between accuracy and latency across various state-of-the-art RAG systems, offering useful context for evaluating ChatRAG’s performance. The CRAG benchmark employs a structured scoring method to assess the quality of responses generated by RAG systems, categorizing each answer as perfect, missing, or incorrect. Following this approach, we implemented a similar labeling process. In our evaluation, we relied on an LLM-as-Judge approach (Gu et al., 2025) to validate the correctness of the final answers, excluding responses such as “I don’t know”. LLM-as-Judge approaches provide efficient ways to evaluate query responses that go beyond simple requests like multiple-choice questions and have proven remarkably effective for many benchmarks, matching even crowdsourced human evaluations (Zheng et al., 2023). We employed DeepSeek-V2 (DeepSeek-AI et al., 2024) as our LLM judge due to its low cost and competent question-answering capabilities. The details of this evaluation and the used prompts are available in our replication package [^13] https://anonymous.4open.science/r/green-rag-techniques-experiment/running/README.md.
@@ 3.6. Data Analysis
After we collected all the data from the experiments, we assessed the normality of each data subset using the Shapiro-Wilk test (Hanusz et al., 2016). Since all datasets in our experiment passed the normality test, we used the t-test (Student, 1908) to evaluate the significance levels. Based on our hypotheses, we used a one-tailed test for energy consumption and a two-tailed test for performance. Since multiple hypotheses were tested simultaneously, we had to guard against the multiple comparisons problem, where the probability of incorrectly rejecting at least one true null hypothesis (Type I error) increases with the number of tests performed (Barnett et al., 2022). To mitigate this, we applied the Holm-Bonferroni correction (Holm, 1979), which adjusts the p-values to control the family-wise error rate. After correction, an adjusted p-value ¡ 0.05 indicates a statistically significant difference, while a value ¿ 0.05 suggests no significant difference. For statistically significant differences, we calculated Cohen’s d (Cohen, 2013) to estimate the effect size. According to Cohen’s guidelines, values between 0.0 and 0.2 indicate a negligible effect, 0.2 to 0.5 a small effect, 0.5 to 0.8 a moderate effect, and 0.8 and above a large effect. Additionally, we calculated mean values for each configuration with a significant effect and then computed the percentage change relative to the baseline (control).
@@ 4.1. Reducing Energy Consumption (RQ1)
This limited effectiveness is likely due to the short prompt lengths and minimal overlap between retrieved documents across queries. To quantify query similarity, we applied TF-IDF vectorization (Rajaraman and Ullman, 2011) combined with cosine similarity. The average similarity score was 0.0294 (standard deviation 0.0437), indicating that most queries differ substantially. As a result, the prefix cache was rarely reused, leading to negligible improvements. In contrast, systems with higher query similarity might benefit more from prefix caching.
@@ 4.3. Trade-off With Accuracy (RQ2b)
In this section, we first examine the baseline’s ability to answer questions accurately and then evaluate the impact of the proposed techniques on system accuracy. Among all 2,606 questions from our benchmark dataset, the baseline system produced, on average, correct answers for 619 questions, resulting in an accuracy of 23.75%. This is in line with the published CRAG benchmark results (Yang et al., 2024), which mention an accuracy of 23.7% for Llama 3.1 8B Instruct. Our baseline system achieved almost identical accuracy, which validates the realism and effectiveness of our base setup, but also of our LLM-as-Judge approach.
@@ 4.4. Combining the Two Best Techniques
Two technique configurations (T1 with the 0.78 threshold and T3 with 384 dimensions) emerged as the most beneficial candidates, improving energy consumption, latency, and even slightly accuracy. To understand whether their benefits compound when applied together, we ran an additional experiment with a system variation that implemented both of them simultaneously. As shown in Fig. 7, the combination of the two led to even greater improvements in energy consumption and latency. Specifically, T1-0.78 alone achieved a 20.7% reduction in energy consumption, while T3-384 alone yielded a 38.1% reduction. Their combination, however, reduced energy consumption by 46.7%. The results for latency were similarly beneficial. Although each individual configuration results in a modest accuracy improvement of 1.7%, their combination did not produce a statistically significant change in accuracy anymore. According to Gu et al. (2021), as the dimensionality of embeddings increases, the distribution of pairwise cosine similarities tends to converge toward a stable distribution with finite variance. In higher-dimensional spaces, cosine similarity scores cluster more tightly around a central value, reducing the variance between similarity scores across different pairs. In practice, high-dimensional embeddings tend to make random vectors more orthogonal, concentrating pairwise cosine similarities near zero with low variance. This shifts the natural baseline similarity upward, e.g., from a slightly negative value in lower dimensions to near zero, which likely requires a threshold recalibration. Nonetheless, since the combination also did not harm accuracy, applying both techniques at once remains a powerful option to improve energy efficiency.
@@ 6. Threats to Validity
We assess potential threats to the validity of our experiment following the framework of Wohlin et al. (Wohlin et al., 2012).
Internal Validity. Chunking is critical for RAG systems, as larger chunks can improve recall but may reduce precision (Pesl et al., 2025). We used the TokenTextSplitter of Spring AI with a chunk size of 320 tokens to balance model compatibility and manageable chunk counts. Despite this, 10 out of 501,916 chunks failed, which could marginally affect accuracy if they contained relevant information.